# Random Forest for Dynamic Risk Prediction of Recurrent Events: A Pseudo-Observation Approach

### v2

@@ 3 Random Forest for Uncensored Longitudinal Outcomes: A Review
Random forests for survival data can be evaluated using Harrell’s C-statistic (also called “C-index”), a non-parametric measure of concordance (Harrell and others, 1982). Briefly, the C-statistic compares the predicted risk scores or fit values for any two observations with the actual time-to-event for those two observations. The proportion of observations that are correctly ordered is reported as the C-index. Any model which perfectly orders patients based on some risk score has a C-index of 1, while a model with no predictive power has C-index equal to 0.50.
In the next section we review pseudo-observation methodology that can convert the censored longitudinal data developed in Section 2 into a version that can be used with the above algorithm.
4 Construction of Pseudo-Observations
Pseudo-observations are a variation of jack-knife methodology applicable to censored time-to-event outcomes (Andersen, Klein and Rosthøj, 2003). This technical device has been used successfully for a single time-to-event and recurrent events in manuscripts such as: Andrei and Murray (2007); Hengelbrock and others (2016); Xia, Murray and Tayob (2020); Zhao and others (2020). Essentially a pseudo-observation of a censored outcome, once created, can be used in regression contexts as if it was an uncensored outcome. In this section, we will first describe the mechanics of how to create a pseudo-observation for each potentially censored time-to-first event in a follow-up window of length $\tau$ starting at $t$. We will then briefly describe properties of pseudo-observations in the context of the random forest described in Section 3.
In this manuscript, our parameter of interest is survival probability, but we recognize that pseudo-observation is a technique that can be applied to a wide variety of other targets, including $\tau$-restricted mean survival time and hazard estimation. To construct pseudo-observations to be used with our random forest algorithm, we return to the censored longitudinal data structure described in Section 2. We calculate two estimates of the probability of recurrent-event-free in a follow-up window of length $\tau$ beginning at check-in time $t$: $\widehat{P}(X_{i}(t)\geq\tau)$ and $\widehat{P}^{(-i)}(X_{i}(t)\geq\tau)$ via the Kaplan-Meier estimator without conditioning on any covariate information for an individual $i=1,2,...,n(t)$ at risk at time $t$, where $\widehat{P}^{(-i)}(X_{i}(t)\geq\tau)$ is estimated excluding individual $i$. The pseudo-observation for individual $i$ at time $t$ is defined by:
$\widehat{S}^{\tau}_{i}(t)=n(t)\widehat{P}(X_{i}(t)\geq\tau)-(n(t)-1)\widehat{P}^{(-i)}(X_{i}(t)\geq\tau).$
To gain intuition behind the use of pseudo-observations, consider the simplest case where there is no censoring. Then, algebraically, $\widehat{P}(T_{i}(t)\geq\tau)=\frac{\sum_{j=1}^{n(t)}I(T_{j}(t)\geq\tau)}{n(t)}$ and $\widehat{P}^{(-i)}(T_{i}(t)\geq\tau)=\frac{\sum_{j\neq i,j=1}^{n(t)}I(T_{j}(t)\geq\tau)}{n(t)-1}$. Here $\widehat{S}_{i}^{\tau}(t)$ reduces to $I(T_{i}(t)\geq\tau)$, so that $E[\widehat{S}_{i}^{\tau}(t)\mid\bm{Z}_{i}(t)]=E[I(T_{i}(t)\geq\tau)|\bm{Z}_{i}(t)]=P(T_{i}(t)\geq\tau|\bm{Z}_{i}(t))$. This property is approximately held when we use Kaplan-Meier estimates for $\widehat{P}(T_{i}(t)\geq\tau)$ and $\widehat{P}^{(-i)}(T_{i}(t)\geq\tau)$ in the independently right-censored case. Andersen, Klein and Rosthøj (2003) argued that this property enables consistent estimation by directly regressing the pseudo-observations upon covariates. When pseudo-observations of probability of time-to-first-recurrent-events are applied to random forests in Section 3 for prediction, we call the algorithm “RFRE.PO” (short for “Random Forest for Recurrent Events based on Pseudo-Observations”); we depict the analysis pipeline in Figure 2.
Figure 2: A flowchart depicting the analysis pipeline of RFRE.PO. Labels above arrows correspond to the section which describes the corresponding step in greater detail. An analyst begins with data recorded as recurrent event times before constructing an appropriate $\mathcal{T}$, and transforming data into the censored longitudinal framework. Upon creating such a structure, the analyst can select the preferred quantity for calculating pseudo-observations, which are ultimately fed to a historical random forest.
@@ 5 Simulation Study
As mentioned earlier in Section 1, random forest algorithms are known for nonparametrically capturing complex relationships between a large number of predictors and outcomes of interest, where these complex relationships might involve higher-order interactions, nonlinearity, and colinearity between predictors. In this section, we demonstrate the ability of RFRE.PO to predict $P(T_{i}(t)\geq\tau\mid\bm{Z}_{i}(t))$ in a scenario where the true relationship between $\bm{Z}_{i}(t)$ and $I(T_{i}(t)\geq\tau)$ is many times more complicated than a modeler who tends to depend on main effect terms would anticipate. For comparison, we also describe results from using a perfectly specified Xia, Murray and Tayob (XMT) model (Xia, Murray and Tayob, 2020) (Model A); and the XMT model that only uses main effect terms for the $\bm{Z}_{i}(t)$ in its regression model (Model B). Note that in Xia, Murray and Tayob (2020), the XMT model was for modeling the restricted mean rather than $\widehat{S}^{\tau}_{i}(t)$; however, their methods can be extended to our case. We did not include highly colinear predictors that would cause modified XMT model convergence issues in our simulations, although random forest algorithms handle colinear predictors well.
@@
In our experience, one of the most important predictors of future recurrent events is past event rate history. This is particularly important in settings where $T_{i}(t)$, $t\in\mathcal{T}_{i}$, are highly correlated within individual $i$. Studies do not always have access to recurrent event history at follow-up time $t=t_{0}$, although this information grows and is available for later follow-up windows. The goals of our simulation study are (1) to verify high performance of our random forest algorithm for prediction of $P(T_{i}(t)\geq\tau|\bm{Z}_{i}(t))$ in the recurrent event setting and (2) to demonstrate the importance of event rate history in modeling $P(T_{i}(t)\geq\tau|\bm{Z}_{i}(t))$ with suggestions on how this can be done. As part of goal (2), for each modeling strategy (RFRE.PO, Model A, Model B), we consider three different settings corresponding to availability of historical event rate information: (1) none, where there is no history covariate available (2) full, where the history covariate is available for each individual $i$ in each follow-up window starting at $t\in\mathcal{T}_{i}$ or (3) partial, where the history covariate is only recorded for windows beginning at $t>t_{0}$.
We now describe the details of how data were simulated for our study. The true relationship between recurrent events ($T_{i,1}<T_{i,2}<\ldots$) and patient characteristics (continuous covariates $Z_{1}$, $Z_{2}$, $Z_{4}$, and categorical covariates $Z_{3}$, $Z_{5}$, $Z_{6}$, $Z_{7}$) is designed to be intentionally complex. We generate the recurrent events by simulating and transforming possibly dependent gap-times by Gaussian copula. The recurrent event gap-times marginally follow an exponential distribution with hazard:
$\lambda_{i}=\exp\{Z_{i2}\sin(Z_{i1}/Z_{i6})-Z_{i3}I(Z_{i2}>2)+Z_{i3}I(Z_{i2}\leq 2)+Z_{i1}Z_{i6}+Z_{i2}^{2}Z_{i4}\}.$
@@
We simulate $Z_{1}\sim N(0,5)$, $Z_{2}\sim N(2,0.8)$, $Z_{3}+2\sim\text{Poisson}(4)$, $Z_{4}-0.1\sim\text{Beta}(7,1)$, $Z_{5}\in\{0,1,2\}$ with probability $\big{(}\frac{2}{3},\frac{1}{6},\frac{1}{6}\big{)}$, $Z_{6}\in\{-5,-2,2,3\}$ with probability $\big{(}\frac{1}{10},\frac{1}{3},\frac{11}{30},\frac{1}{5}\big{)},$ and $Z_{7}\in\{0,1,2,3\}$ with equal probability. Individuals with simulated hazards outside of the range of $\frac{8}{15}$ and $15$ years are removed from the sample and replaced. We induce correlation among exponentially distributed gap times for each subject $i$ by utilizing the Gaussian copula method as described by Xia, Murray and Tayob (2020), for $\rho=0,0.3,0.6$ or $0.9$. The resultant recurrent event times are converted to a censored longitudinal data format as described in Section 2. Censoring is independently generated from an exponential distribution so that 0%, $23\%$, $45\%$, or $63\%$ of observations are censored before the end of the study.
[…]
For each of the models described above, as well as RFRE.PO, Table 1 lists the mean and empirical standard deviation of the index of concordance of the algorithms discussed and Figure 2 shows the distribution of the index of concordance of the algorithms discussed. Increasing levels of correlation between recurrent event times are listed from left to right. As one moves down the table and figure, censoring levels increase.
For all levels of censoring, correlation and historical covariate use, excluding the highest correlation level with historical covariate use, Model A (the ad hoc nonlinear modified XMT model) has the best C-index with varying degrees of superiority to Model B (the main effects modified XMT model) and RFRE.PO method. The RFRE.PO approach outperforms Model B in all settings, except high correlation with use of covariate history, in which Model B competes with both RFRE.PO and Model A. This suggests that in highly correlated settings, patient history is more important than baseline covariate information.
In the absence of correlation, use of $H_{p}(t)$ in Model B weakens predictive performance, whereas predictive performance is similar for the Model B with $H_{f}(t)$ versus without use of $H_{f}(t)$. The RFRE.PO procedure is able to make large gains in predictive ability from use of $H_{p}(t)$ and $H_{f}(t)$ in the same case with no correlation. Generally, use of $H_{f}(t)$ gives superior predictive performance to use of $H_{p}(t)$ in settings with correlation between recurrent event times. Our summary of the RFRE.PO results is that any use of covariate history in settings with and without correlated recurrent event outcomes is quite important and assists in recovering information from the data as censoring increases.
Table 1: Simulated C-statistics listed as Mean (ESD) for varying prediction algorithms, censoring percentages (none is 0%, light is 23%, moderate is 45%, and heavy is 63%), and gap-time correlations ($\rho=0$, $0.3$, $0.6$, $0.9$). 300 replicates were performed with $n=500$ for each replicate.
${\rho=0}$ ${\rho=0.3}$ ${\rho=0.6}$ ${\rho=0.9}$
No Censoring
Model A 0.677 (0.006) 0.672 (0.007) 0.664 (0.008) 0.651 (0.010)
RFRE.PO 0.598 (0.014) 0.589 (0.015) 0.576 (0.016) 0.558 (0.016)
Model B 0.530 (0.007) 0.531 (0.009) 0.531 (0.010) 0.529 (0.011)
Model A with $H_{p}(t)$ 0.675 (0.006) 0.674 (0.007) 0.684 (0.007) 0.710 (0.007)
RFRE.PO with $H_{p}(t)$ 0.629 (0.007) 0.651 (0.007) 0.683 (0.008) 0.710 (0.008)
Model B with $H_{p}(t)$ 0.489 (0.023) 0.582 (0.019) 0.666 (0.011) 0.724 (0.009)
Model A with $H_{f}(t)$ 0.676 (0.006) 0.677 (0.006) 0.691 (0.011) 0.712 (0.010)
RFRE.PO with $H_{f}(t)$ 0.634 (0.008) 0.658 (0.007) 0.686 (0.009) 0.710 (0.007)
Model B with $H_{f}(t)$ 0.519 (0.035) 0.645 (0.035) 0.685 (0.017) 0.731 (0.010)
Light Censoring
Model A 0.676 (0.006) 0.672 (0.008) 0.665 (0.009) 0.650 (0.011)
RFRE.PO 0.594 (0.014) 0.585 (0.014) 0.574 (0.016) 0.557 (0.016)
Model B 0.530 (0.009) 0.531 (0.009) 0.532 (0.010) 0.530 (0.011)
Model A with $H_{p}(t)$ 0.676 (0.007) 0.673 (0.008) 0.681 (0.007) 0.704 (0.008)
RFRE.PO with $H_{p}(t)$ 0.628 (0.008) 0.649 (0.008) 0.680 (0.006) 0.707 (0.007)
Model B with $H_{p}(t)$ 0.491 (0.025) 0.574 (0.019) 0.659 (0.013) 0.717 (0.011)
Model A with $H_{f}(t)$ 0.675 (0.006) 0.677 (0.007) 0.689 (0.010) 0.705 (0.010)
RFRE.PO with $H_{f}(t)$ 0.631 (0.008) 0.655 (0.008) 0.684 (0.008) 0.707 (0.008)
Model B with $H_{f}(t)$ 0.523 (0.037) 0.642 (0.017) 0.680 (0.019) 0.724 (0.011)
Moderate Censoring
Model A 0.676 (0.007) 0.672 (0.007) 0.664 (0.009) 0.649 (0.012)
RFRE.PO 0.591 (0.014) 0.584 (0.015) 0.570 (0.015) 0.555 (0.017)
Model B 0.531 (0.009) 0.532 (0.010) 0.531 (0.010) 0.530 (0.012)
Model A with $H_{p}(t)$ 0.675 (0.007) 0.672 (0.008) 0.679 (0.008) 0.701 (0.008)
RFRE.PO with $H_{p}(t)$ 0.623 (0.009) 0.644 (0.008) 0.676 (0.007) 0.705 (0.007)
Model B with $H_{p}(t)$ 0.492 (0.023) 0.565 (0.021) 0.649 (0.015) 0.712 (0.012)
Model A with $H_{f}(t)$ 0.675 (0.007) 0.676 (0.007) 0.686 (0.011) 0.700 (0.011)
RFRE.PO with $H_{f}(t)$ 0.629 (0.009) 0.653 (0.008) 0.680 (0.009) 0.704 (0.008)
Model B with $H_{f}(t)$ 0.528 (0.034) 0.639 (0.016) 0.674 (0.021) 0.718 (0.012)
Heavy Censoring
Model A 0.676 (0.007) 0.672 (0.009) 0.664 (0.010) 0.649 (0.011)
RFRE.PO 0.586 (0.017) 0.577 (0.016) 0.568 (0.018) 0.551 (0.017)
Model B 0.532 (0.009) 0.532 (0.009) 0.533 (0.012) 0.530 (0.012)
Model A with $H_{p}(t)$ 0.674 (0.007) 0.672 (0.009) 0.676 (0.009) 0.695 (0.009)
RFRE.PO with $H_{p}(t)$ 0.618 (0.010) 0.640 (0.009) 0.671 (0.008) 0.700 (0.007)
Model B with $H_{p}(t)$ 0.492 (0.025) 0.554 (0.022) 0.640 (0.016) 0.703 (0.013)
Model A with $H_{f}(t)$ 0.675 (0.007) 0.675 (0.008) 0.684 (0.011) 0.695 (0.012)
RFRE.PO with $H_{f}(t)$ 0.625 (0.010) 0.650 (0.009) 0.676 (0.009) 0.700 (0.009)
Model B with $H_{f}(t)$ 0.531 (0.032) 0.636 (0.016) 0.668 (0.024) 0.711 (0.014)
Figure 3: A violin plot of the distributions of the C-statistic of the models under consideration in simulation. Measures of center and spread can be found in Table 1. Within a box, historical covariate information increases from left to right. For partial history information, the distribution for the average C-statistic from $m=10$ imputed data sets is displayed.

### v3

@@ 6 Evaluation Metrics
Both the m-XMT models and RFRE.PO may be evaluated using Harrell’s C-statistic (also called C-index or the index of concordance), a non-parametric measure of concordance (Harrell and others, 1982). Briefly, the C-statistic compares the predicted risk scores or fit values for any two observations with the actual time-to-event for those two observations. The proportion of observations that are correctly ordered is reported as the C-statistic. Event pairs where censoring obscures the correct ordering of events within the pair are not included in the C-statistic calculation. Any model that perfectly orders patients based on some risk score has a C-statistic of 1, while a model with no predictive power has C-statistic equal to 0.50. C-statistics for events in the censored longitudinal data structure given in Section 2 can be reported across all events or within $\tau$-duration follow-up windows starting at $t$. In addition to C-statistics, overall and time-dependent mean squared error (MSE) estimates are used to assess model calibration as described in Supplemental Materials Section LABEL:s:supplementMSE.
@@ 7 Simulation Study
As mentioned in Section 1, random forest algorithms are known for nonparametrically capturing complex relationships between a large number of predictors and the outcome of interest, where these complex relationships might involve higher-order interactions, nonlinearity, and collinearity between predictors. In this section, we evaluate the ability of RFRE.PO to predict $P\{T_{i}(t)\geq\tau\mid\bm{Z}_{i}(t)\}$ in a scenario where the true relationship between $\bm{Z}_{i}(t)$ and $I\{T_{i}(t)\geq\tau\}$ is many times more complicated than a modeler who tends to use main effect terms would anticipate. For comparison, we also describe results from using a perfectly specified modified XMT model, labeled m-XMT (True), and a modified XMT model that only uses main effect terms for $\bm{Z}_{i}(t)$ in equation (11), labeled m-XMT (Main Effects).
@@
The goals of our simulation study are (1) to verify high performance of our random forest algorithm for prediction of $P\{T_{i}(t)\geq\tau|\bm{Z}_{i}(t)\}$ in the recurrent event setting and (2) to demonstrate the importance of event rate history in modeling $P\{T_{i}(t)\geq\tau|\bm{Z}_{i}(t)\}$. As part of goal (2), for each modeling strategy [RFRE.PO, m-XMT (True), m-XMT (Main Effects)], we consider three different settings corresponding to availability of historical event rate information as described in Section 2: (1) none, (2) $H_{p}^{(1)}(t)$, or (3) $H_{f}^{(1)}(t)$.
@@
We now describe the details of how data were simulated. The true relationship between recurrent events ($T_{i,1}<T_{i,2}<\ldots$) and patient characteristics (continuous covariates $Z_{1}$, $Z_{2}$, $Z_{4}$, and categorical covariates $Z_{3}$, $Z_{5}$, $Z_{6}$, $Z_{7}$) is designed to be intentionally complex. Recurrent event gap-times marginally follow an exponential distribution with hazard: $\lambda_{i}=\exp\{Z_{i2}\sin(Z_{i1}/Z_{i6})+(-1)^{I(Z_{i2}>2)}Z_{i3}+Z_{i1}Z_{i6}+Z_{i2}^{2}Z_{i4}\}.$ We simulate $Z_{1},Z_{2}\sim\text{MVN}(\mathbf{\mu},\Sigma)$, where $\mathbf{\mu}=(0,2)^{\top}$ and $\Sigma=\begin{pmatrix}5&0.7\\ 0.7&.8\end{pmatrix}$, $Z_{3}+2\sim\text{Poisson}(4)$, $Z_{4}-0.1\sim\text{Beta}(7,1)$, $Z_{5}\in\{0,1,2\}$ with probability $\big{(}\frac{2}{3},\frac{1}{6},\frac{1}{6}\big{)}$, $Z_{6}\in\{-5,-2,2,3\}$ with probability $\big{(}\frac{1}{10},\frac{1}{3},\frac{11}{30},\frac{1}{5}\big{)},$ and $Z_{7}\in\{0,1,2,3\}$ with equal probability. The sign in front of $Z_{3}$ depends on the value of $Z_{2}$ which is correlated with $Z_{1}$. $Z_{5}$ and $Z_{7}$ are noise covariates. Individuals with simulated hazards corresponding to mean times-to-recurrent event outside of the range of $\frac{1}{15}$ and $\frac{15}{8}$ years are removed from the sample and replaced. We induce correlation among exponentially distributed gap times for each subject $i$ by utilizing the Gaussian copula method as described by Xia, Murray and Tayob (2020), for $\rho=0,0.3,0.6$ or $0.9$. The resultant recurrent event times are converted to a censored longitudinal data format as described in Section 2. Censoring is independently generated from an exponential distribution so that 0%, $23\%$, $45\%$, or $63\%$ of individuals are censored.
The censored longitudinal data structure assumes $\mathcal{T}=\{0,\frac{1}{12},\frac{1}{6},...2\}$ years, and $\tau=\frac{1}{6}$ year. For individual $i$, $\mathcal{T}_{i}$ is the subset of follow-up window start-times in $\mathcal{T}$ where individual $i$ remains at risk for recurrent events. Both the m-XMT (True) model and the m-XMT (Main Effects) estimate $\text{logit}\{E(S^{\tau}_{i}(t))\}$ assuming the relationship $\text{logit}\{E(S^{\tau}_{i}(t))\}=\mathbf{\beta}^{\top}\bm{Z}^{*}_{i}(t)$. For m-XMT (True), the ad hoc nonlinear model, $\bm{Z}^{*}_{i}(t)=\{Z_{i2}\sin(Z_{i1}/Z_{i6}),$ $(-1)^{I(Z_{i2}>2)}Z_{i3},$ $Z_{i1}Z_{i6},$ $Z_{i2}^{2}Z_{i4}\}^{\top}$. The m-XMT (Main Effects) model has covariate vector $\bm{Z}^{*}_{i}(t)=\{Z_{i1},$ $Z_{i2},$ $Z_{i3},$ $Z_{i4},$ $Z_{i5},$ $Z_{i6},$ $Z_{i7},t\}^{\top}$. Covariate histories are included as additional patient characteristics in $\bm{Z}^{*}_{i}(t)$ for full and partial history availability settings. For full history availability settings this predictor takes the form $H_{f,i}^{(1)}(t)$ as given in Equation (1) with $\mathcal{T}^{*}_{i}(t)=\{\frac{-1}{12}\}\cup\{t_{k}\in\mathcal{T}_{i}:t_{k}+\tau^{*}<t\}$, and $\tau^{*}=\frac{1}{12}$. For partial history availability settings, imputes for missing event histories prior to $t=0$ follow recommendations in Section 2 with $\tilde{T}_{i,j}\sim$Exp$(\lambda_{i})$, where $\lambda_{i}\sim$Unif$(\frac{8}{15},15)$. Partial history covariates are added to $\bm{Z}^{*}_{i}(t)$ based on Equation (2), with the same $\mathcal{T}^{*}_{i}(t)$ and $\tau^{*}$ as $H_{f}^{(1)}(t)$. All finite sample performance results are based on 500 simulated datasets of n=500 participants per scenario.
Figures 3A and 3B display simulated model fit for (1) the modified XMT model using main effects terms only {m-XMT (Main Effects), shown in purple}, (2) our random forest algorithm for recurrent events (RFRE.PO, shown in green) and (3) the true modified XMT model {m-XMT (True), shown in yellow}. Panel (A) displays violin plots of C-statistics (indices of concordance) seen in simulation with corresponding measures of center and spread shown in Table LABEL:tab:simResultTable. For partial history information, average C-statistics from $m=10$ imputed data sets are shown. Within a box, historical covariate information increases from left to right. Panel (B) expands on the scenario boxed in red from panel (A), i.e., the low censoring setting with 0.3 correlation. Here violin plot C-statistic distributions are time-dependent, reflecting model fit for $\tau=\frac{1}{6}$ duration follow-up windows starting at times, $t$, marked on the horizontal axis. The quality of historical information increases from top to bottom. The m-XMT (True) model performance provides a benchmark for assessing performance of the proposed (truth-blind) dynamic prediction methods, the m-XMT (Main Effects) model and the RFRE.PO algorithm.
Panel A results show that when there is no historical covariate used for dynamic prediction, the m-XMT (True) model has the best C-statistic performance, as expected, for all levels of censoring and correlation, followed by the RFRE.PO algorithm and then the m-XMT (Main Effects) model. As correlation between recurrent event times increases, use of either partial or full history covariates in these latter two dynamic prediction algorithms brings C-statistic performance closer to that of the m-XMT (True) model. When using full history covariates, RFRE.PO outperforms the m-XMT (Main Effects) model for $\rho=0.3$ and performs similarly for $\rho=0.6$. When using partial history covariates, RFRE.PO outperforms m-XMT (Main Effects) for both $\rho=0.3$ and $\rho=0.6$. For $\rho=0.9$, use of either partial or full history covariates in the main effects model slightly outperforms both the RFRE.PO algorithm and the m-XMT (True) model.
Generally, the inclusion of full history covariates is more effective than that of partial history covariates, though the benefit of having full history covariates available is most strongly seen in the lowest two correlation settings when applying the m-XMT (Main Effects) model. In the highest correlation setting, partial and full covariate history algorithms have very similar performance. In the absence of correlation between recurrent event times, use of history covariates shows modest, if any, improvement for the RFRE.PO algorithm C-statistic whereas C-statistic performance for the m-XMT (Main Effects) model deteriorates noticeably.
Time-dependent C-statistic performance highlighted in Panel B shows how quickly algorithms using partial and full covariate history information improve and stabilize over the course of follow-up. Stabilization happens quickly, with little change in time-dependent follow-up windows after $t=5$. Because covariate history variables improve with increasing $t$, there is an induced interaction between these predictors and $t$, particularly when the algorithms are using the partial history covariates. This type of interaction is naturally assessed by the RFRE.PO approach.
Together, Figures 3A and 3B indicate that in highly correlated settings, (1) patient history is an essential predictor to include in dynamic prediction algorithms and (2) the RFRE.PO algorithm seems to outperform the m-XMT (Main Effects) algorithm in all but the very highest correlation setting where history covariates are used for dynamic prediction. Additional simulation results are available in Supplemental Materials. Section LABEL:s:supplementMSEInterp summarizes mean squared error results with corresponding Table LABEL:tab:simMSETable, while Section LABEL:s:supplementImportance describes the ability of these algorithms to distinguish signal covariates ($Z_{1},Z_{2},Z_{3},Z_{4},Z_{6}$) versus noise covariates ($Z_{5},Z_{7}$) in Tables LABEL:tab:m-XMTSig and LABEL:tab:checkPower.